OpenStudy (anonymous):

A star of radius 1.3 × 105 km rotates about its axis with a period of 36 days. The star undergoes a supernova explosion, whereby its core collapses into a neutron star of radius 11 km. Estimate the period of the neutron star (assume the mass remains constant). Answer in units of s

OpenStudy (anonymous):

You can use conservation of angular momentum and treat each body as a sphere \[L = I\omega\] \[L_I = L_f\] \[ I\omega_{Big \ Star} = I \omega_{Neutron \ Star}\] \[I_{sphere} = \frac{2}{5}mr^2\] solve for the angular velocity of the neutron start, then convert to the period \[\omega = 2\pi f=\frac{2 \pi}{T}\]

OpenStudy (anonymous):

what does L and I stand for?

OpenStudy (anonymous):

L is angular momentum and I is the moment of inertia ^_^ Have you used those before? Can you find the angular velocity of the star given its period of rotation?

OpenStudy (anonymous):

I've never used this before :/ maybe that's why it's so hard ha

OpenStudy (anonymous):

^_^ You've done normal conservation of momentum before, right?

OpenStudy (anonymous):

\[ p_i = p_f \\ \ \\ mv_i = mv_f\] Conservation is angular momentum is exactly like that, exact we use something called a cross product \[\textbf r_i \times m \textbf v_i = \textbf r_f \times m \textbf v_f\] Luckily for this problem, that simplifies down to \[I_i\omega_i = I_f \omega_f\] The moment of inertia, I, of an object is kinda like a term that shows "an objects tendency to resist changes in motion" - which is exactly what "mass" does in a normal momentum problem. In a table you can look up that the moment of inertia of a sphere is \[I_{sphere} = \frac{2}{5}mr^2\] and the angular velocity, omega, of the original sun is given by 2 pi divided by its period of rotation (which is saying that it goes 2 pi radians, or one complete revolution, in T seconds.) \[\omega_{sun} = \frac{2 \pi}{T_{sun}}\] \[I_{sun}\omega_{sun} = I_{neutron} \omega_{neutron}\] \[\Big(\cancel{\frac{2}{5}}\cancel{m}r^2_{sun} \Big)\omega_{sun} = \Big(\cancel{\frac{2}{5}}\cancel{m}r^2_{neutron} \omega_{neutron}\] Notice that writing out both terms from the definitions of moment of inertia lets the masses cancel out ^_^ Can you see where to go from here?

OpenStudy (anonymous):

Yes thanks for your help :)